# Deep learning-driven likelihood-free parameter inference for 21-cm forest observations

### v1

@@ Results
To confirm the reliability of our parameter estimates, we conducted the Kolmogorov-Smirnov (KS) test 39. The KS test is a non-parametric test that compares the cumulative distribution function (CDF) of the sample data to a reference distribution. By computing the CDF of the percentiles of the true values of each parameter derived from the marginal 1D posterior distribution, we can evaluate the model’s performance. For each 21-cm forest 1D power spectrum and the corresponding confidence interval of the posterior distribution, the percentile values of the injected parameters are computed. Ideally, the CDF for each parameter should be close to the diagonal, as the percentiles should be evenly distributed between 0 and 1. We measured the CDF of the true parameters falling within specific confidence levels as presented in Supplementary Fig. 0.5. The results demonstrated the CDF curves within the $3\sigma$ confidence bands, with $p$-values greater than 0.05 for each parameter. A KS test $p$-value greater than 0.05 indicates that there is no significant difference between the CDF of the sample data and the reference distribution, confirming that the observed deviations are within the expected range due to random sampling variability. These findings affirm the validity and accuracy of our inference method within the expected error ranges, providing confidence in its application for cosmological studies. The KS test results support the reliability of our approach in accurately estimating $m_{\rm WDM}$ and $T_{\rm K}$, ensuring that our method is theoretically sound and practically reliable.
@@ Discussion
Observing the 21-cm forest signals during the EoR provides a valuable opportunity to gain insight into DM properties and the heating history in the first billion years of the universe. However, due to the lack of explicit relation between the physical parameters and the observables, parameter inference has been impeded by the limited parameter space covered by computationally expensive simulations and the infeasible requirement on computational resources. Moreover, the non-Gaussian characteristic of the 21-cm forest signal complicates the evaluation of the likelihood function, resulting in biased inference of parameters. In this paper, we introduce a deep learning-based approach that enables model-independent inference of $T_{\rm K}$ and $m_{\rm WDM}$ with a small number of simulations. Notably, this approach could extend to non-Gaussian scenarios, allowing for parameter inference using arbitrary 21-cm forest signals.
Furthermore, the deep learning-based parameter inference method presented in this work demonstrates innovation and effectiveness not only in astronomy and cosmology but also in handling complex and unbalanced datasets 40. For example, in biomedical research, where sample scarcity hampers understanding of rare diseases 41, our advanced data analysis technology can significantly enhance data analysis capabilities and drive scientific discovery. With ongoing algorithm optimization and advancements in computing power, this method is poised to play a pivotal role in advancing interdisciplinary research.
In our work, we employ the multi-scale modeling approach from Ref. 11 to simulate the 21-cm forest signal. We use 21cmFAST 42 to model the ionization field 43, 44 and density field 45, 46, 47, 48 on large scales, identifying neutral regions to simulate the 21-cm signal, with the density fields serving as initial conditions for small-scale simulations 10, 49, 50, 51, 52. The main contributors to the 21-cm signal are minihalos, modeled using the conditional halo mass function 45, 46 to estimate the number and distribution of low-mass minihalos, with a minimum halo mass of $10^{6}M_{\odot}$, corresponding to the filter mass 53 at the redshift of interest. Halo density profiles are calculated using the Navarro-Frenk-White (NFW) profile 54 within the virial radius and the infall model profile 55 outside it. The infall model, suitable for high-density areas with numerous halos, tends to overestimate the density field, requiring normalization of the field to match the initial local density of 21cmFAST outputs. The normalization technique of Ref. 11, which involves multiplying each voxel by a factor, alters the amplitude of the 1D power spectrum of the 21-cm forest and leads to a crossover in the 1D power spectrum across different DM models (See Fig. 3 in Ref. 11). To address this, we employ a normalization method that does not affect the amplitude of the 1D power spectrum, adjusting the density field by adding or subtracting a number. Furthermore, to maintain the probability density distribution of the density field, we adjust the lower limit of halo mass to $10^{5}M_{\odot}$ in the simulation, corresponding to the Jeans mass 56 at the redshift of interest, which also mitigates anomalies in the 1D power spectrum on large scales. Figure 6 in Ref. 11 shows that the degeneracy direction between $m_{\rm WDM}$ and $T_{\rm K}$ is negatively correlated. As we previously discussed in Supplementary Fig. 0.5, different scales affect the degeneracy direction of the two parameters differently. In this work, the improvements we made have altered the relationship of the 1D power spectrum on large scales for different DM models. Consequently, in Fig. 6, the degeneracy direction between $m_{\rm WDM}$ and $T_{\rm K}$ becomes positively correlated. Meanwhile, using SKA2-LOW with an integration time of 100 h is insufficient to observe the 21-cm forest signal. As a result, we increased the integration time to 200 h.
@@
For data analysis, we calculate the two-point correlation function of the 21-cm forest and convert it to the $k$-space, resulting in the 1D power spectrum along the line of sight. The choice stems from the faintness of the 21-cm forest signal, making a power spectrum analysis more beneficial. Nonetheless, other studies 26 suggest alternative methods that could more effectively compress data and yield more precise parameter constraints. Future research will explore these methods for nearly lossless parameter inference, enhancing the effectiveness of constraints. Additionally, we plan to incorporate factors like metal absorption lines 57 coupling and foreground to create a more accurate dataset.
@@
Overall, our study demonstrates that deep learning-based methods, particularly NFs, offer a promising approach to analyzing the 21-cm forest signal. These methods address the challenges of simulation data scarcity and complexity, providing robust and accurate parameter inference. As the SKA era approaches, these innovative techniques will be invaluable for extracting maximum scientific information from 21-cm observations, contributing to a deeper understanding of the thermal history of the universe and the properties of DM.
@@ Methods
where $\nu_{0}$ = 1420.4 MHz is the rest frame frequency of 21-cm photons, $T_{\mathrm{S}}$ is the spin temperature of HI gas, $\tau_{\nu_{0}}$ is the 21-cm optical depth. In terms of average gas properties within each voxel, the 21-cm optical depth can be written as 3, 58, 59
@@
where $\delta(\hat{\mathbf{s}},z)$, $x_{\mathrm{HI}}(\hat{\mathbf{s}},z)$ and $H(z)$ are the gas overdensity, the neutral fraction of hydrogen and the Hubble parameter, respectively. ${\mathrm{d}v_{\|}/\mathrm{d}r_{\|}}$ is the inherent velocity projected along the line of sight gradient and we ignore this term in the simulation.
For the spin temperature, we assume that it is coupled to $T_{\rm K}$ ($T_{\rm S}\approx T_{\rm K}$). The gas temperature in each voxel is set by the thermal history of the early universe and its proximity to halos. In this study, our focus is on neutral regions, where X-ray penetration primarily heats the gas, and subsequently affects the 21-cm signals. When we consider an unheated IGM ($f_{\rm X}$ = 0), the temperature of gas outside the virial radius is mainly determined by the cosmic expansion. For the brightness temperature of the background radiation $T_{\gamma}$, our assumption is the same as Ref. 11, which is related to the flux density of the background source $S_{150}$. In this work, we consider 10 background sources with $S_{150}$ = 10 mJy at $z$ = 9 and perform a total of 100 measurements on 10 comoving Mpc length segments along the line of sight in a neutral patch.
@@
Using a hybrid multi-scale approach, we simulate the 21-cm forest during the EoR. We employ the 21cmFAST to model large-scale structures in a (1 Gpc)3 box with $500^{3}$ grids, initialized by DM properties. Further, each grid with a length of 2 Mpc is subdivided into $500^{3}$ voxels. These voxels are populated with halos based on initial densities and the conditional halo mass function influenced by the $m_{\rm WDM}$. Density within each voxel follows either the NFW model or the infall model 55, depending on proximity to the nearest halo. The model is as presented in Ref. 11. Throughout the study, we adopted a set of cosmological parameters consistent with Planck 2018 results 60: $\Omega_{\mathrm{m}}=0.3153$, $\Omega_{\mathrm{b}}h^{2}=0.02236$, $\Omega_{\Lambda}=0.6847$, $h=0.6736$ and $\sigma_{8}=0.8111$. $\Omega_{\Lambda}$ and $\sigma_{8}$ are the dark energy density parameter and the matter fluctuation amplitude, respectively.
@@
Observational uncertainties in the 21-cm forest include thermal noise, sample variance, contaminating spectral structure of foreground sources in the chromatic side lobes, and bandpass calibration errors 11. Regarding the error consideration, we only consider the thermal noise of the interferometer array and the sample variance in the 1D power spectrum measurement as Ref. 11. The sample variance is generated by the simulation process and does not need to be added separately. When measuring individual absorption lines directly, the noise flux density averaged over the two polarizations can be expressed as 61
@@
where $\lambda_{z}$ is the observation wavelength, $\Omega=\pi(\theta/2)^{2}$ is the solid angle of the telescope beam, where $\theta$ = $1.22\lambda_{z}/D$ is the angular resolution, $D$ is the longest baseline of a radio telescope or array. For SKA1-LOW, we use $A_{\mathrm{eff}}/T_{\mathrm{sys}}$ = 800 $\rm{m^{2}/K}$ 62. For SKA2-LOW, we use $A_{\mathrm{eff}}/T_{\mathrm{sys}}$ = 4,000 $\rm{m^{2}/K}$ 63. For these two arrays, we assume $D$ = 65 km, and the integration time $\delta t$ is 100 h and 200 h, respectively.
In Supplementary Fig. 0.5, we calculated the brightness temperature for different cases. Direct measurement of individual absorption lines is easily hampered by early X-ray heating. To improve the sensitivity of detecting the 21-cm forest signal and reveal the clustering properties of the absorption lines to distinguish the effect on the 21-cm signal between the heating effect and WDM model, we followed the algorithm in Ref. 11 and calculated 1D power spectrum of brightness temperature under the assumption of high redshift background sources 64. The brightness temperature $\delta T_{\mathrm{b}}(\hat{\mathbf{s}},\nu)$ as a function of the observation frequency $\nu$ can be equivalently used as the line-of-sight distance $r_{z}$, $T^{\prime}_{\mathrm{b}}(\hat{\mathbf{s}},r_{z})$ to represent. The Fourier transform of $T^{\prime}_{\mathrm{b}}(\hat{\mathbf{s}},r_{z})$ is
@@
Here $1/{\Delta r_{z}}$ is the normalization factor, where ${\Delta r_{z}}$ is the line of sight length considered. To reveal the small-scale structures we are interested in, we select ${\Delta r_{z}}$ = 10 comoving Mpc neutral patches and calculate the 1D power spectrum of the 10 comoving Mpc segment along the line of sight. For a reasonable number $\mathcal{O}$ (10) high-$z$ background sources, the expected value of the 1D power spectrum is obtained by incoherently averaging over 100 neutral patches on the line of sight penetrating various environments, i.e. $P\left(k_{\|}\right)\equiv\left\langle P\left(\hat{\mathbf{s}},k_{\|}\right)\right\rangle$. In the rest of this work, we abbreviate $k_{\|}$ to $k$ because here we are always interested in $k$ along the line of sight. We calculated the 1D power spectrum as presented in Supplementary Fig. 0.5 of the 21-cm forest corresponding to different conditions. The dotted line indicates the thermal noise of SKA2-LOW when the integration time is 200 hr. It can be seen that using the 1D power spectrum can improve the sensitivity of detecting 21-cm forest signals.
@@
We finally obtained the simulated correspondence between parameters ($m_{\rm WDM}$, $T_{\rm K}$) and the 1D power spectrum in the presence of thermal noise. Specifically, we simulated 7 simulation boxes with DM particle masses ranging from 3 keV to 9 keV at equal intervals (1 keV). For the gas temperature, we consider two situations. When the heating effect is weak ($f_{\rm X}\approx 0.1$), the IGM will be heated to about 60 K at $z=9$. Therefore, we choose equal intervals (5 K) between 40 K and 80 K. When the heating effect is strong ($f_{\rm X}\approx 1$), the IGM will be heated to about 600 K at $z=9$. Therefore, we choose equal intervals (100 K) between 400 K and 800 K. In addition, we also generated some data sets of other parameters for validation and test sets. To avoid the influence of the edge effect, we only select the synthetic spectrum of $30\times 300$ in each simulation box to calculate the 1D power spectrum. It should be noted that this method is not limited to the above parameter range and is also applicable to other parameter ranges.
@@
$\tilde{P}(k)=2\frac{\log P(k)-\log P_{\max}(k)}{\log P_{\max}(k)-\log P_{\min}(k)}-1,$ (8)
where $P(k)$ represents the value of the 1D power spectrum at a specific wavenumber $k$ after binning, $\log P_{\max}(k)$ is the logarithm of the maximum 1D power spectrum value at $k$, and $\log P_{\min}(k)$ is the logarithm of the minimum 1D power spectrum value at $k$. This approach helps balance the contributions of each scale and improves the stability and efficiency of network training. The data set generated and processed by the above method as presented in Supplementary Fig. 0.5, which not only provides a solid foundation for training and testing our deep learning model but also ensures the reliability and scientific significance of the research results.

### v2

@@ Results
To assess the performance of the INF in deriving cosmological parameters, we evaluate two key aspects: the reliability of the inferred parameters and the precision of the parameter estimates. First, we need to confirm the reliability of the INF inference parameters. We conducted the Kolmogorov-Smirnov (KS) test 43. The KS test is a non-parametric test that compares the cumulative distribution function (CDF) of the sample data to a reference distribution. By computing the CDF of the percentiles of the true values of each parameter derived from the marginal 1D posterior distribution, we can evaluate the model’s performance. For each 21-cm forest 1D power spectrum and the corresponding confidence interval of the posterior distribution, the percentile values of the injected parameters are computed. Ideally, the CDF for each parameter should be close to the diagonal, as the percentiles should be evenly distributed between 0 and 1. We measured the CDF of the true parameters falling within specific confidence levels as presented in Fig. 5. The results demonstrated the CDF curves within the $3\sigma$ confidence bands, with $p$-values greater than $0.05$ for each parameter. A KS test $p$-value greater than $0.05$ indicates that there is no significant difference between the CDF of the sample data and the reference distribution, confirming that the observed deviations are within the expected range due to random sampling variability. These findings affirm the validity and accuracy of our inference method within the expected error ranges, providing confidence in its application for cosmological research.
Next, we determine the precision of the parameters inferred by INF. Similar to findings in Refs. 30, 25, we encounter challenges due to limited prior parameter bounds, as probabilities outside the predefined range cannot be estimated. However, this can be solved by expanding the coverage of the simulated parameter space. Additionally, using reciprocal parameter values in transformations could improve the inference framework’s applicability, making it possible to avoid simulating to infinity. Within the central regions of the parameter space, our analysis reveals distinct patterns in the relative error $\varepsilon$ of parameter inference for $m_{\rm WDM}$ and $T_{\rm K}$ under various heating levels, as presented in Supplementary Fig. 0.8. Here, the relative error, defined as the precision, is given by $\varepsilon(\xi)=\sigma(\xi)/\xi$, where $\xi$ represents the true parameter ($m_{\rm WDM}$ or $T_{\rm K}$). Although SKA2-LOW is more sensitive, the higher error is due to the fact that our discussion for SKA1-LOW focused on cases with weaker heating effects, while increasing the heating effect suppresses the 21-cm forest signal, leading to larger inference errors. We find that $T_{\rm K}$ significantly affects the precision of $m_{\rm WDM}$ estimates. Conversely, $m_{\rm WDM}$ has a relatively small impact on the precision of $T_{\rm K}$ estimates. This is because $T_{\rm K}$ influences the 1D power spectrum values across all scales, while $m_{\rm WDM}$ primarily affects the 1D power spectrum values on small scales, as shown in Supplementary Fig. 0.8. Therefore, as presented in Supplementary Fig. 0.8, $T_{\rm K}$ can achieve high precision by relying solely on large-scale information, almost unaffected by $m_{\rm WDM}$. On the other hand, $m_{\rm WDM}$ mainly relies on small-scale and medium-scale information, which is influenced by $T_{\rm K}$. Therefore, the 1D power spectrum of different scales will break the degeneracy, resulting in a significant improvement in the precision of parameter estimates. This dependence highlights the interplay between these parameters and emphasizes the ability of our network to effectively constrain each parameter. Generating $2^{16}$ posterior samples with INF takes only $32.2$ ms, showcasing its remarkable efficiency, far surpassing the speed of GNF combined with Bayesian inference.
For low heating level conditions, we utilized SKA1-LOW with an integration time of $100$ h to examine the constraint results of posterior parameter inference, assuming $m_{\rm WDM}=6$ keV and $T_{\rm K}=60$ K. We compared the precision of the posterior inference results from the INF with the range predicted by the Fisher matrix as presented in Fig. 6. The INF method successfully recovered the posterior distribution of parameters, with the true values lying within the estimated $1\sigma$ confidence region. This demonstrates precision comparable to the predictions of the Fisher matrix. The final average constraint result of the 1D power spectra of the 21-cm forest for all simulations are $m_{\rm WDM}=6.28_{-1.59}^{+1.61}$ keV and $T_{\rm K}=60.17_{-6.66}^{+6.16}$ K. This confirms the reliability of the INF method in achieving precise parameter inference under low heating conditions.
Under high heating level conditions, using SKA2-LOW with an integration time of $200$ h ($m_{\rm WDM}=6$ keV, $T_{\rm K}=600$ K), we observed significantly different results. Although both the Fisher matrix and INF methods presented consistent degeneracy directions, the Fisher matrix approach failed to adequately constrain the parameters due to the non-Gaussian likelihood function of the 1D power spectrum of the 21-cm forest. This highlights the effectiveness of likelihood-free inference methods like the INF in such scenarios. The final average constraint results of the 1D power spectra for all simulations are $m_{\rm WDM}=6.47_{-1.88}^{+2.04}$ keV and $T_{\rm K}=627.95_{-91.07}^{+94.18}$ K. These results demonstrate that, while the Fisher matrix approach is inadequate under a non-Gaussian likelihood function, the INF method remains effective.
@@ Discussion
Observing the 21-cm forest signals during the EoR provides a valuable opportunity to gain insight into DM properties and the heating history in the first billion years of the universe. However, due to the lack of explicit relation between the physical parameters and the observables, parameter inference has been impeded by the limited parameter space covered by computationally expensive simulations and the infeasible requirement on computational resources. Moreover, the non-Gaussian likelihood function of the 21-cm forest signal complicates the evaluation of the likelihood function, resulting in biased inference of parameters. In this work, we introduce a deep learning-based approach that does not rely on traditional analytical models, enabling inference of $T_{\rm{K}}$ and $m_{\rm WDM}$ with a small number of simulations, unlike analytical models that establish an explicit relationship between $m_{\rm WDM}$, $T_{\rm{K}}$ during the reionization period, and the statistical characteristics of the 21 cm signal. Furthermore, this approach could extend to other non-Gaussian scenarios, allowing for parameter inference using arbitrary 21-cm forest signals.
@@
For data analysis, we calculate the two-point correlation function of the 21-cm forest and convert it to the $k$-space, resulting in the 1D power spectrum along the line of sight. The choice stems from the faintness of the 21-cm forest signal, making a power spectrum analysis more beneficial. Nonetheless, other studies 27 suggest alternative methods that could more effectively compress data and yield more precise parameter constraints. Future research will explore these methods for nearly lossless parameter inference, enhancing the effectiveness of constraints. Additionally, we plan to incorporate factors like metal absorption lines 44 coupling and foreground to create a more accurate dataset. To improve the astrophysical modeling, the inclusion of spatial temperature non-uniformity and X-ray source distribution variations 45, 46 is also worth considering. Additionally, the impact of including additional parameters that influence the 21-cm forest warrants investigation, where methods like Latin hypercube sampling (LHS) could potentially enhance sampling efficiency 47, 48.
@@
Overall, our research demonstrates that deep learning-based methods, particularly NFs, offer a promising approach to analyzing the 21-cm forest signal. These methods address the challenges of simulation data scarcity and complexity, providing robust and accurate parameter inference. In addition, the INF method can assist the 21-cm forest in distinguishing between the CDM and WDM models. Given the computationally intensive nature of CDM simulations, we conservatively use the simulated upper limit of $9$ keV as a stand-in for CDM in our analysis. In Supplementary Fig. 0.8, we assess the discriminative power of the INF method by comparing the confidence intervals with the upper limit ($9$ keV) of the parameter space for the simulated $m_{\rm WDM}$ under different $m_{\rm WDM}$ and $T_{\rm K}$. The results indicate that the INF method can distinguish the 1D power spectrum of $m_{\rm WDM}$ values less than 6 keV from that of the CDM model at a $2\sigma$ confidence level. more, the deep learning-based parameter inference method demonstrated here not only advances astronomy and cosmology but also shows potential for handling complex and irregular datasets 49. In fields like biomedical research, where sample scarcity hinders the study of rare diseases 50, our method can significantly improve data analysis capabilities. As the SKA era approaches, these innovative techniques will be invaluable for extracting maximum scientific information from 21-cm observations, contributing to a deeper understanding of the thermal history of the universe and the properties of DM.
@@ Methods
where $\nu_{0}=$ 1420.4 MHz is the rest frame frequency of 21-cm photons, $T_{\mathrm{S}}$ is the spin temperature of HI gas, $\tau_{\nu_{0}}$ is the 21-cm optical depth. In terms of average gas properties within each voxel, the 21-cm optical depth can be written as 3, 51, 52
@@
where $\delta(\hat{\mathbf{s}},z)$, $x_{\mathrm{HI}}(\hat{\mathbf{s}},z)$ and $H(z)$ are the gas overdensity, the neutral fraction of hydrogen and the Hubble parameter, respectively. ${\mathrm{d}v_{\|}/\mathrm{d}r_{\|}}$ is the inherent velocity projected along the line of sight gradient and we ignore this term in the simulation. Additionally, doppler shifts due to any peculiar velocities are also ignored.
For the spin temperature, we assume that it is coupled to $T_{\rm K}$ ($T_{\rm S}\approx T_{\rm K}$), due to the Wouthuysen-Field effect 53, 54. The gas temperature in each voxel is set by the thermal history of the early universe and its proximity to halos. Here, our focus is on neutral regions, where X-ray penetration primarily heats the gas, and subsequently affects the 21-cm signals. When we consider an unheated IGM ($f_{\rm X}=0$), the temperature of gas outside the virial radius is mainly determined by the cosmic expansion. The 21-cm forest signal is more prominent on smaller scales, and thus our work prioritizes the small-scale features. As described in Ref. 45, while X-ray heating introduces temperature variations on large scales due to the spatial distribution of X-ray sources, the long mean free path of X-ray photons ensures that, on scales of $\sim 10$ Mpc, the background temperature can be reasonably approximated as uniform. Additionally, we also account for the temperature inhomogeneities within halos and in their surrounding regions: inside halos, the temperature is set to the virial temperature, while outside halos, it depends on X-ray heating. This approach ensures that the influence of small-scale variations is included in our modeling. In addition, our work primarily considers the upper limit of halos with a virial temperature of $10^{4}{\rm~{}K}$. Below this threshold, it is difficult for stars and galaxies to form within the halo. Thus, these halos do not contain X-ray sources, allowing us to assume a relatively uniform distribution of X-ray sources, leading to uniform heating of IGM. While spatial variations in heating may exist in reality 46, this approximation serves as a starting point for this work, which prioritizes the development of an inference method that can be extended to incorporate more complex physical scenarios in future work. For the brightness temperature of the background radiation $T_{\gamma}$, our assumption is the same as that in Ref. 11, which is related to the flux density of the background source $S_{150}$.
@@
We employ the multi-scale modeling approach proposed in Ref. 11 to simulate the 21-cm forest signal. Specifically, we use 21cmFAST 55 to model the ionization field 56, 57 and density field 58, 59, 60, 61 on large scales in a ($1$ Gpc)3 box with $500^{3}$ grids. Subsequently, we identify neutral regions within the large-scale box to simulate the 21-cm signal, where each grid with a length of $2$ Mpc is further divided into $500^{3}$ voxels by small-scale simulations. The initial conditions for the grids are derived from large-scale density field simulations. Since low-mass halos are the main contributors to the 21-cm forest signal 10, 62, 63, 64, 65, we calculate their number and distribution within small-scale grid based on the conditional mass function. The density of each voxel is determined by its nearest halo, following the Navarro-Frenk-White profile 66 within the virial radius and the infall model profile 67 outside of it. Throughout this work, we adopted a set of cosmological parameters consistent with Planck 2018 results 68: $\Omega_{\mathrm{m}}=0.3153$, $\Omega_{\mathrm{b}}h^{2}=0.02236$, $\Omega_{\Lambda}=0.6847$, $h=0.6736$ and $\sigma_{8}=0.8111$. Here, $\Omega_{\mathrm{m}}$ is the matter density, $\Omega_{\mathrm{b}}$ is the baryonic density, $\Omega_{\Lambda}$ is the dark energy density, $h$ is the dimensionless hubble parameter, and $\sigma_{8}$ is the matter fluctuation amplitude.
@@
Observational uncertainties in the 21-cm forest include thermal noise, sample variance, contaminating spectral structure of foreground sources in the chromatic side lobes, and bandpass calibration errors 11. Regarding the error consideration, we only consider the thermal noise of the interferometer array and the sample variance in the 1D power spectrum measurement as in Ref. 11. The sample variance is generated by the simulation process and does not need to be added separately. When measuring individual absorption lines directly, the noise flux density averaged over the two polarizations can be expressed as 69
@@
where $\lambda_{z}$ is the observation wavelength, $\Omega=\pi(\theta/2)^{2}$ is the solid angle of the telescope beam, where $\theta$ = $1.22\lambda_{z}/D$ is the angular resolution, $D$ is the longest baseline of a radio telescope or array. For SKA1-LOW, we use $A_{\mathrm{eff}}/T_{\mathrm{sys}}=$ $800$ $\rm{m^{2}/K}$ 39, 38. For SKA2-LOW, we use $A_{\mathrm{eff}}/T_{\mathrm{sys}}=$ $4000$ $\rm{m^{2}/K}$ 39, 40. For these two arrays, we assume $D=$ $65$ km, and the integration time $\delta t$ is $100$ h and $200$ h, respectively.
In Supplementary Fig. 0.8, we calculated the brightness temperature for different cases. Direct measurement of individual absorption lines is easily hampered by early X-ray heating. To improve the sensitivity of detecting the 21-cm forest signal and reveal the clustering properties of the absorption lines to distinguish the effect on the 21-cm signal between the heating effect and WDM model, we followed the algorithm in Ref. 11 and calculated 1D power spectrum of brightness temperature under the assumption of high redshift background sources 70. The brightness temperature $\delta T_{\mathrm{b}}(\hat{\mathbf{s}},\nu)$ as a function of the observation frequency $\nu$ can be equivalently used to represent the line-of-sight distance $r_{z}$ as $T^{\prime}_{\mathrm{b}}(\hat{\mathbf{s}},r_{z})$. The Fourier transform of $T^{\prime}_{\mathrm{b}}(\hat{\mathbf{s}},r_{z})$ is
@@
Here $1/{\Delta r_{z}}$ is the normalization factor, where ${\Delta r_{z}}$ is the line of sight length considered. For a reasonable number $\mathcal{O}$ (10) high-$z$ background sources 12, the expected value of the 1D power spectrum is obtained by incoherently averaging over neutral patches on the line of sight penetrating various environments, i.e. $P\left(k_{\|}\right)\equiv\left\langle P\left(\hat{\mathbf{s}},k_{\|}\right)\right\rangle$. The background sources are assumed to be located outside the other end of the simulation volume, with lines of sight randomly directed through the simulation, each crossing multiple neutral regions. This setup ensures a realistic representation of the spatial distribution of neutral patches and the effects of varying local density. In the rest of this work, we abbreviate $k_{\|}$ to $k$ because here we are always interested in $k$ along the line of sight. We calculated the 1D power spectrum as presented in Supplementary Fig. 0.8 of the 21-cm forest corresponding to different conditions. The dotted line indicates the thermal noise of SKA2-LOW when the integration time is $200$ h. It can be seen that utilizing the 1D power spectrum, the 21-cm forest signals can be detected.
Within the large-scale simulation box, we randomly selected 10 neutral hydrogen regions corresponding to 10 background radio sources at redshift $z=9$ with flux densities $S_{150}=10~{}{\rm mJy}$. Each region contains 5 adjacent grid cells along the line of sight for detailed small-scale modeling. By tessellating these 2 Mpc grid units, we established a composite grid system spanning 2 Mpc $\times$ 2 Mpc $\times$ 10 Mpc, containing $500\times 500\times 2500$ voxels (i.e., $500\times 500$ independent line-of-sight paths, each penetrating a 10 Mpc-depth structure). This configuration preserves the 10 Mpc-scale correlations derived from the 21cmFAST simulations at large scales while introducing negligible artifacts in the 1D small-scale power spectrum along the line of sight. To mitigate edge effects caused by excluding external DM halos in our simulation, we implemented boundary truncation in two transverse dimensions: removing peripheral layers from the original $500\times 500$ line array perpendicular to the line-of-sight direction, retaining only the central $300\times 300=90000$ valid line-of-sight paths. For each background source, we assumed that 10 distinct 10 Mpc neutral hydrogen segments could be extracted. Given potential intervening ionized bubbles between neutral segments, we adopted a conservative estimate of 200 Mpc total line-of-sight length, equivalent to a redshift span $\Delta z\approx 0.8$ — a range consistent with practical observational constraints. To effectively suppress thermal noise, we simulated 90000 power spectra per background source. For each source, these power spectra were partitioned into 9000 groups, each containing 10 line-of-sight paths. All 10 paths within a group originate from identical simulation parameters, thereby representing 10 neutral hydrogen segments in a single background source spectrum. Finally, through incoherent averaging of 100 power spectra (10 sources $\times$ 10 segments), we generated 9000 synthetic 1D power spectra.
However, the infall model, suitable for high-density areas with numerous halos, tends to overestimate the density field, requiring normalization of the field to match the initial local density of 21cmFAST outputs. The normalization technique of Ref. 11, which involves multiplying each voxel by a factor, alters the amplitude of the 1D power spectrum of the 21-cm forest and leads to a crossover in the 1D power spectrum across different DM models (See Fig. $3$ in Ref. 11). To address this, we employ a normalization method that does not affect the amplitude of the 1D power spectrum, adjusting the density field by adding or subtracting a number. Furthermore, to maintain the probability density distribution of the density field, we adjust the lower limit of halo mass to $10^{5}M_{\odot}$ in the simulation, corresponding to the Jeans mass 71 at the redshift of interest, which also mitigates anomalies in the 1D power spectrum on large scales. These adjustments to the density field and halo mass lower limit modify the structure of the 1D power spectrum at different scales, thereby changing the direction of parameter degeneracies. Figure $6$ in Ref. 11 shows that the degeneracy direction between $m_{\rm WDM}$ and $T_{\rm K}$ is negatively correlated. As we previously discussed in Supplementary Fig. 0.8, different scales affect the degeneracy direction of the two parameters differently. In this work, the improvements we made have altered the relationship of the 1D power spectrum on large scales for different DM models. Consequently, in Fig. 6, the degeneracy direction between $m_{\rm WDM}$ and $T_{\rm K}$ becomes positively correlated. Meanwhile, using SKA2-LOW with an integration time of $100$ h is insufficient to observe the 21-cm forest signal. As a result, we increased the integration time to $200$ h.
@@
We finally obtained the simulated correspondence between parameters ($m_{\rm WDM}$, $T_{\rm K}$) and the 1D power spectrum in the presence of thermal noise. Specifically, we simulated $7$ simulation regions with DM particle masses ranging from $3$ keV to $9$ keV at equal intervals ($1$ keV). For the gas temperature, we consider two situations. When the heating effect is weak ($f_{\rm X}\approx 0.1$), the IGM will be heated to about $60$ K at $z=9$. Therefore, we choose equal intervals ($5$ K) between $40$ K and $80$ K. When the heating effect is strong ($f_{\rm X}\approx 1$), the IGM will be heated to about $600$ K at $z=9$. Therefore, we choose equal intervals ($100$ K) between $400$ K and $800$ K. For each set of parameters, out of 9000 power spectra, we take 300 as the testing set and use the remaining 8700 for training. In addition, we also generated some data sets of other parameters for validation and testing sets as presented in Supplementary Fig. 0.8. It should be noted that this method is not limited to the above parameter range and is also applicable to other parameter ranges. Although LHS 47, 48 could reduce the number of sampling points per dimension, the computational cost of varying $T_{\rm{K}}$ is relatively small once $m_{\rm WDM}$ is fixed, making LHS less beneficial in this context. Specifically, for the density field simulation, when the mass is 9 keV, it can take up to $1317$ h on $64$ cores (Intel(R) Xeon(R) Gold 6271C CPU @ 2.60GHz). In contrast, the simulation of temperature field can be completed within 1 min. Moreover, given the correlation between $m_{\rm WDM}$ and $T_{\rm{K}}$, this approach would be less effective. For future studies involving more parameters, exploring higher-dimensional spaces and considering methods like LHS to improve efficiency are worthwhile directions for further investigation. However, in studies of the 21-cm forest, the dominant factors influencing the signal arise from DM properties governing the density field and heating effects modulating the temperature field. Cosmological parameters exhibit minimal observational impact during the EoR, while astrophysical parameters linked to reionization processes remain subject to significant uncertainties. We therefore prioritize these two pivotal parameters that critically shape the 21-cm forest signal, as they encapsulate the dominant physical mechanisms relevant to near-future investigations.
@@
$\tilde{P}(k)=2\frac{\log P(k)-\log P_{\min}(k)}{\log P_{\max}(k)-\log P_{\min}(k)}-1,$ (8)
where $P(k)$ represents the value of the 1D power spectrum at a specific wavenumber $k$ after binning, $\log P_{\max}(k)$ is the logarithm of the maximum 1D power spectrum value at $k$, and $\log P_{\min}(k)$ is the logarithm of the minimum 1D power spectrum value at $k$. This approach helps balance the contributions of each scale and improves the stability and efficiency of network training. The data set generated and processed by the above method are presented in Supplementary Fig. 0.8, which not only provides a solid foundation for training and testing our deep learning model but also ensures the reliability and scientific significance of the research results. Additionally, we conducted a test on two completely random points, and the results, as shown in Supplementary Fig. 0.8, validate the model.